# Debunking with Dialogue? Exploring AI-Generated Counterspeech to Challenge Conspiracy Theories

### v1

@@ Abstract
Counterspeech is a key strategy against harmful online content, but scaling expert-driven efforts is challenging. Large Language Models (LLMs) present a potential solution, though their use in countering conspiracy theories is under-researched. Unlike for hate speech, no datasets exist that pair conspiracy theory comments with expert-crafted counterspeech. We address this gap by evaluating the ability of GPT-4o, Llama 3, and Mistral to effectively apply counterspeech strategies derived from psychological research provided through structured prompts. Our results show that the models often generate generic, repetitive, or superficial results. Additionally, they over-acknowledge fear and frequently hallucinate facts, sources, or figures, making their prompt-based use in practical applications problematic.
@@
Mareike Lisker  and Christina Gottschalk  and Helena Mihaljević University of Applied Sciences (HTW) Berlin, Germany
@@ 1 Introduction
Manual moderation for comprehensive counterspeech is challenging due to high volumes and the rapidly evolving social media landscape Bonaldi et al. (2024a). In this context, supporting NGOs by automatically generating counterspeech that requires little post-processing becomes increasingly vital Mun et al. (2024). Large Language Models (LLMs) could potentially assist Chung et al. (2019). While research has explored using LLMs for counterspeech against HS, there is a notable gap in addressing CTs. While CTs can incite hatred and often perpetuate antisemitic or racist narratives Samantha Hay (2020), they typically differ linguistically from HS.
@@
Our findings show that all three models struggled to produce CS with substantial depth, often defaulting to superficial statements and avoiding narrative storytelling as a strategy. Responses frequently contained generic and repetitive phrases, resulting in limited linguistic variety. Notably, approximately 10% of outputs contained hallucinations which were often difficult to spot. Differences between the models were low and mostly not significant. These findings suggest that current LLMs are not yet effective for the generation of CS to CTs in a zero-shot setting, underscoring the need for further research on grounding LLMs in relevant domain knowledge.
@@ 2 Related Work
Research on automated CS generation has gained traction, but has primarily focused on HS. To our knowledge, only Costello et al. (2024) have examined the use of LLMs to generate CS for CTs. In lab experiments, participants engaged in multi-turn dialogues with GPT-4 Turbo, resulting in a statistically significant reduction in self-reported CT belief. This suggests general potential for CT belief revision through LLM interaction, though in a direct, dialogue-based setting differing from the social media context. We found no publicly available datasets containing examples of effective CS to CTs communicated online. Existing datasets on online CTs (e. g., Langguth et al. (2023); Steffen et al. (2023)) were mainly created for classification tasks Liu et al. (2024); Peskine et al. (2023); Pustet et al. (2024), and are mostly related to COVID-19.
@@
The majority of approaches in CS generation rely on fine-tuning methods Qian et al. (2019); Tekiroglu et al. (2022), using datasets from platforms such as Reddit and Gab Qian et al. (2019), compiled by NGO operators Chung et al. (2019), or generated through models Bonaldi et al. (2022). More recently, zero-shot and few-shot prompting of various GPT-based models has shown promising results for generating CS against HS Chung et al. (2021); Ashida and Komachi (2022); Zheng et al. (2023); Halim et al. (2023). Some studies have explored tailored approaches, such as using type-prompts to specify predefined styles like humor or facts Saha et al. (2024) or personalizing CS based on author profiling (e. g., age and gender) Doğanç and Markov (2023) to fit specific contexts. Across these studies, the newer GPT models demonstrate strong performance.
@@
In addition to quantitative methods, most researchers rely on human annotators for more in-depth evaluations with regard to ‘informativeness’ and ‘offensiveness’ Ashida and Komachi (2022), or ‘suitableness’, and ‘intra-coherence’ Chung et al. (2021), ‘grammaticality’, ‘specificity’ Tekiroglu et al. (2022), or ‘diversity’, ‘relevance’ and ‘language quality’ Zhu and Bhat (2021).
@@ 3 Counterspeech Against Conspiracy Theories
Conspiracy theories are interpretative frameworks attempting to explain events as the covert actions of powerful, malicious groups or individuals pursuing self-interest at public expense Aaronovitch (2010); Byford (2011); Keeley (1999), overlooking actual, more intricate causes Popper and Kiesewetter (2003). They thus usually need actors (e. g., ‘globalists’) with a malicious goal (e. g., destabilizing national economies) through an action or strategy (e. g., manipulating weather) carried out in secrecy Samory and Mitra (2018). On social media, these narratives often appear in fragmented forms, with some components omitted or implied Steffen et al. (2023). While CTs can overlap with HS Baider (2023), they have their own unique characteristics with regard to function and linguistic manifestation Samantha Hay (2020).
@@ 4.1 Dataset
The data was collected using keywords associated with two CT themes, which can be broadly characterized as hate-based and fear-driven (see Table 1). In total, we collected 152 posts (76 per category). All data and methods are available in the project repository. [^1] https://github.com/ai-researcher-anon/debunking_with_dialogue
@@ 4.3 Prompt Design
We opted for zero-shot prompting as no reference CS was available. We follow recommendations by LLM providers and recent research (cf., e. g., Schulhoff et al. (2024)), by including role and style prompting, as well as step-by-step instructions. The instructions primarily reflect four strategies for countering CTs that, according to current literature, are likely to enhance the effectiveness of CS: fact-check based refutation, providing alternative explanations, storytelling of a counter-narrative, and encouraging critical thinking. For the theoretical underpinning of the prompt design, see Appendix A.
@@ 4.4 Counterspeech Annotation Framework
Since comparing model output with ‘correct’ counterspeech through quantitative metrics such as BLEU or METEOR cannot be employed, we rely on manual annotation of model outputs based on the prompt instructions. We annotated a total of 456 responses across 12 criteria, each reflecting a specific aspect of the prompt instructions. The criteria are color-coded to correspond to the respective step in the prompt (see Appendix B) as well as the respective literature they were derived from (see A):
@@
To evaluate the annotation schema, a random sample of 50 responses—presented without any information about the respective models—was annotated independently by two authors. The inter-annotator agreement, measured using Krippendorf’s Alpha, ranged from 0.35 (Conciseness and clarity) to 0.83 (Hate speech condemnation and Appropriate empathy for fear). Disagreements were discussed and resolved collaboratively, and the annotation schema was refined accordingly. The remaining data was split equally and annotated by one author each in close communication, with $\sim\!\!25\%$ of the records discussed jointly. Likert scales were iteratively refined, and prior annotations were retrospectively updated to ensure consistency.
@@ 4.5 Comparison of Models
We quantitatively compared how the models implemented CS strategies based on the annotated Likert scale ratings. Due to the non-normal distribution of the values, we used the non-parametric Friedman test ($\alpha=0.05$) to compare all three models. For significant results, we applied the the post-hoc Wilcoxon Signed-Rank test to identify model pairs with significant differences. Effect sizes were measured using Kendall’s W. Comparisons were limited to records with non-missing values for all models. Bonferroni correction was applied to adjust p-values and control the family-wise error rate.
@@ 5.1 Adherence to Prompt Instructions
The models showed varying adherence to simple prompt instructions, such as listing how hate speech was handled and which strategies were applied (the ‘meta report’), separating this information from the CS using the <XXX> tag, and keeping the CS under 200 tokens. GPT-4o consistently included the meta report, while Llama 3 and Mistral omitted it in 26–28% of outputs (see col. 1 in Table 3). The meta report in GPT-4o and Llama 3 responses was short, accounting for $\sim 1\%$ and $\sim 4\%$ of the output, respectively, while Mistral produced substantially longer explanations ($>20\%$ on average). GPT-4o correctly separated the meta report from the generated CS in over 96% of outputs, enabling easy rule-based parsing. In contrast, Llama 3 responses required manual post-processing in 55% of cases and Mistral failed to use the tag <XXX>, necessitating manual post-processing for all outputs. Mistral reported on HS handling in only 25% of outputs, compared to 56% for GPT-4o and 68% for Llama 3 (see col. 3 in Table 3). Interestingly, all models frequently misinterpreted the requirement to acknowledge fear—when present—as a CS strategy, often labeling it as ‘Strategy 5,’ thus indicating confusion about the prompt instructions.
@@
Overall, acknowledgments were often superficial, frequently appearing in introductory remarks such as, “It’s natural to feel concerned about […]. However, […].” Consequently, average scores were low, from $1.9\pm 0.86$ (Mistral) to $1.95\pm 0.65$ (GPT-4o) and $2.15\pm 0.58$ (Llama 3), with no statistically significant differences. In some cases, acknowledging fear reinforced a conspiratorial mindset, e. g., “It’s natural to feel concerned when leaders address global issues […]” 364.
@@ 5.2 Application of Counterspeech Strategies
In $\sim\!\!10\%$ of the fact-checked responses, models declared true events to be false, hallucinated facts, or produced ‘bullshit’ Frankfurt (2009), such as: “However, it’s important to note that the term ‘globalist’ is often used as a euphemism for ‘those who want to help the world’ or ‘those who believe in international cooperation.”’ 55 Often, the cited sources were untraceable despite sounding convincing e. g., 150, 217, 275. Regularly, though, the models just missed the mark, with a number actually being correct but from another report 302, 154, a number being correct but describing a different context in a report 154, or a quote being partially correct but altogether hallucinated 344.
@@
In terms of named entities, despite the prompt encouraging expert or study citations, CS seldom introduced new individuals beyond those in the input. Except for three historical figures, additional names provided by the models were either hallucinated or tied to incorrect information.
@@
Given that Narr was effectively applied only in 3 out of 456 cases, a second experiment was carried out for exploratory purposes. The prompt was limited to producing a narrative, thus excluding Fact, Alt, and Crit and the handling of fear, while the other steps remained the same. In experiment 2, the revised prompt was applied to the data from the geo-/bio-engineering related CT theme, as the respective messages were less fragmented and thus better suited to produce a counter narrative. It was tested exclusively on GPT-4o, as, overall, it demonstrated the best results and most obedience. One author evaluated the 76 responses, in particular with regard to narrative storytelling and conciseness and clarity.
In this setting, GPT-4o generated narratives in $\sim\!\!60\%$ of the responses. Often, the narratives fit the context but failed to address the core of the CT, yielding a mean Likert score of 2.82 (±0.78). The story of Katalin Krekó was referenced three times and represented a particularly compelling narrative. In other cases, scientists’ names were mentioned without forming a strong narrative 2.404442’ refers to experiment 2, and ‘40’ is the ID of the response., or generic references were made to ‘scientists’ who were presumably intended to be central figures in a story 2.70. One figure, “Jessica the pilot”, was entirely hallucinated 2.48, while the other persons were real individuals including scientists such as James Hansen, Shi Zhengli, Kizzmekia Corbett, or Jennifer Francis, politicians such as Aneurin Bevan, and historical figures such as Alexander Graham Bell or Ada Lovelace. Additionally, public figures such as science communicator Neil deGrasse Tyson, activist Greta Thunberg, and entrepreneur Hamdi Ulukaya were included. Overall, $\sim\!\!40\%$ of the figures were female.
@@ 6 Conclusion and Discussion
All models primarily attempted to counter CTs using fact-checking (Fact), alternative explanations (Alt), and encouragment of critical thinking (Crit). The use of narratives (Narr), despite being explicitly defined and exemplified in the prompt, was exceedingly rare. This strategy being the only one in experiment 2 resulted in a $\sim\!\!60\%$ realization, though not necessarily effective. This suggests that Narr is currently the least suitable strategy for our NGO employee scenario and warrants specific research attention if used in experiments.
Outputs often lacked depth and specificity, with generic responses and overuse of boilerplate text, aligning with recent findings indicating that LLMs tend to generate generic replies, e. g. by simply denouncing a statement Mun et al. (2023), an effect reinforced through safety guardrails of models Bonaldi et al. (2024b). leading to mediocre Likert-scale scores for conciseness and clarity, and strategy effectiveness.
@@
Despite explicit instructions in the prompt, all models hallucinated factual information, including sources, publications, and quotes, sometimes in subtle ways that required detailed scrutiny to identify. Logical fallacies were also observed, further questioning the reliability of current language models for CS in practical settings. These issues, coupled with low diversity in responses, indicate that significant manual post-processing would be necessary to avoid sounding robotic or unreliable in real-world applications while providing factually correct and engaging CS.
@@ Appendix B Prompt
The LLMs were instructed by the following system prompt, while the 152 messages from the social media platform X were part of the prompt. You are a trained expert in generating counterspeech to conspiracy theory comments on X (formerly Twitter). Your goal is to persuade the audience who is undecided about the conspiracy theory not to believe it. Follow these steps while generating counterspeech: Step 1: Maintain Conciseness and Clarity

### v2

@@ Abstract
Counterspeech is a key strategy against harmful online content, but scaling expert-driven efforts is challenging. Large Language Models (LLMs) present a potential solution, though their use in countering conspiracy theories is under-researched. Unlike for hate speech, no datasets exist that pair conspiracy theory comments with expert-crafted counterspeech. We address this gap by evaluating the ability of GPT-4o, Llama 3, and Mistral to effectively apply counterspeech strategies derived from psychological research provided through structured prompts. Our results show that the models often generate generic, repetitive, or superficial results. Additionally, they over-acknowledge fear and frequently make up facts, sources, or figures, making their prompt-based use in practical applications problematic.
@@
Mareike Lisker  and Christina Gottschalk  and Helena Mihaljević University of Applied Sciences (HTW) Berlin, Germany {lisker,mihalje}@htw-berlin.de
@@ 1 Introduction
Manual moderation for comprehensive counterspeech is challenging due to high volumes and the rapidly evolving social media landscape Bonaldi et al. (2024a). In this context, supporting NGOs by automatically generating counterspeech that requires little post-processing becomes increasingly vital Mun et al. (2024). Large Language Models (LLMs) could potentially assist Chung et al. (2019). While research has explored using LLMs for counterspeech against HS, there is a notable gap in addressing CTs. While CTs can incite hatred and often perpetuate antisemitic or racist narratives Hay (2020), they typically differ linguistically from HS.
@@
Our findings show that all three models struggled to produce CS with substantial depth, often defaulting to superficial statements and avoiding narrative storytelling as a strategy. Responses frequently contained generic and repetitive phrases, resulting in limited linguistic variety. Notably, approximately 10% of outputs contained confabulations which were often difficult to spot. Differences between the models were low and mostly not significant. These findings suggest that current LLMs are not yet effective for the generation of CS to CTs in a zero-shot setting, underscoring the need for further research on grounding LLMs in relevant domain knowledge.
@@ 2 Related Work
Research on automated CS generation has gained traction, but has primarily focused on HS. To our knowledge, only Costello et al. (2024) have examined the use of LLMs to generate CS for CTs. In lab experiments, participants engaged in multi-turn dialogues with GPT-4 Turbo, resulting in a statistically significant reduction in self-reported CT belief. This suggests general potential for CT belief revision through LLM interaction, though in a direct, dialogue-based setting differing from the social media context.
We found no publicly available datasets containing examples of effective CS to CTs communicated online. Existing datasets on online CTs (e. g., Langguth et al. (2023); Steffen et al. (2023)) were mainly created for classification tasks Liu et al. (2024b); Peskine et al. (2023); Pustet et al. (2024), and are mostly related to COVID-19.
@@
The majority of approaches in CS generation rely on fine-tuning methods Qian et al. (2019); Tekiroğlu et al. (2022), using datasets from platforms such as Reddit and Gab Qian et al. (2019), compiled by NGO operators Chung et al. (2019), or generated through models Bonaldi et al. (2022). More recently, zero-shot and few-shot prompting of various GPT-based models has shown promising results for generating CS against HS Chung et al. (2021); Ashida and Komachi (2022); Zheng et al. (2023); Halim et al. (2023). Some studies have explored tailored approaches, such as using type-prompts to specify predefined styles like humor or facts Saha et al. (2024) or personalizing CS based on author profiling (e. g., age and gender) Doğanç and Markov (2023) to fit specific contexts. Across these studies, the newer GPT models demonstrate strong performance.
@@
In addition to quantitative methods, most researchers rely on human annotators for more in-depth evaluations with regard to ‘informativeness’ and ‘offensiveness’ Ashida and Komachi (2022), or ‘suitableness’, and ‘intra-coherence’ Chung et al. (2021), ‘grammaticality’, ‘specificity’ Tekiroğlu et al. (2022), or ‘diversity’, ‘relevance’ and ‘language quality’ Zhu and Bhat (2021).
@@ 3 Counterspeech Against Conspiracy Theories
Conspiracy theories are interpretative frameworks attempting to explain events as the covert actions of powerful, malicious groups or individuals pursuing self-interest at public expense Aaronovitch (2010); Byford (2011); Keeley (1999), overlooking actual, more intricate causes Popper and Kiesewetter (2003). They thus usually need actors (e. g., ‘globalists’) with a malicious goal (e. g., destabilizing national economies) through an action or strategy (e. g., manipulating weather) carried out in secrecy Samory and Mitra (2018). On social media, these narratives often appear in fragmented forms, with some components omitted or implied Steffen et al. (2023). While CTs can overlap with HS Baider (2023), they have their own unique characteristics with regard to function and linguistic manifestation Hay (2020).
@@ 4.1 Dataset
The data was collected using keywords associated with two current and rather well-known CT themes, which can be broadly characterized as hate-based and fear-driven, namely theories revolving around ‘deep state’, ‘NWO’ and ‘globalists’ on the hate-based side, and those relating to ‘geo- and bioengineering’ on the fear-based side. The two sets of keywords (see Table 1) were iteratively enriched following an exploration of conspiratorial posts and hashtags on X. Note that we did not aim to create an exhaustive set of keywords, as we are mainly interested in identifying texts that can be clearly categorized as conspiratorial by an LLM. In total, we collected 152 posts (76 per category). All data and methods are available in the project repository. [^1] https://github.com/HTW-Social-Data-Science/Debunking_with_Dialogue
@@ 4.3 Prompt Design
We opted for zero-shot prompting, including role and style prompting, as well as step-by-step instructions, as recommended by LLM providers and recent research (cf., e. g., Schulhoff et al. (2024)). The instructions primarily reflect four strategies for countering CTs that, according to current literature, are likely to enhance the effectiveness of CS: fact-check based refutation (Fact), providing alternative explanations (Alt), storytelling of a counter-narrative (Narr), and encouraging critical thinking (Crit). While Fact involves correcting misinformation through fact-based rebuttals, Alt emphasizes providing more holistic alternative explanations. Narr uses narrative techniques to engage the audience emotionally, whereas Crit aims at revealing the manipulative techniques used in conspiracy theories and encouraging reflective questioning.
In the following, we illustrate the different counterspeech strategies by sample answers to an example CT. For the theoretical underpinning of the prompt design, see Appendix A.
CT: “COVID-19 was created as a bio-weapon in a lab to control the population.”
Fact: “Independent scientific bodies, including the World Health Organization and international virology experts, have found no evidence that COVID-19 was engineered or intentionally released, as noted in a 2020 article by Kristian G. Andersen in Nature Medicine. Genetic analyses indicate natural evolution, and similar coronaviruses found in animals, especially bats, support a zoonotic origin as stated in Wikipedia.”
Alt: “Viruses like COVID-19 have emerged before through natural processes. Scientists have found that the virus shares much of its genetic code with coronaviruses found in bats and other animals, suggesting it likely jumped from animals to humans—something that has happened before with SARS and MERS. In a world where people and animals are increasingly in close contact, and global travel is constant, it’s more likely that this virus emerged through natural transmission rather than being engineered.”
Narr: “In the 1990s, Dr. Lam discovered a virus that had jumped from bats to humans in Malaysia. When COVID-19 emerged, some suspected a lab origin, but Dr. Lam knew that the genetic code of the virus told a different story: it closely matched viruses found in bats, showing a natural origin. He explained that as humans encroach on wildlife habitats, viruses can spill over, and that this is what likely happened with COVID-19.”
Crit: “Mhh, but if COVID-19 was a bio-weapon designed for control, why would it spread uncontrollably across all countries, including those allegedly behind it? Would a government really create something that harms its own economy, overwhelms its hospitals, and puts its own leaders at risk?”
@@ 4.4 Counterspeech Annotation Framework
Since comparing model output with ‘correct’ counterspeech through quantitative metrics such as BLEU or METEOR cannot be employed, we rely on manual annotation of model outputs based on the prompt instructions.
We annotated a total of 456 responses across 12 criteria, each reflecting a specific aspect of the prompt instructions. The criteria are color-coded to correspond to the respective step in the prompt (see Appendix B) as well as the respective literature they were derived from (see A):
@@
[…]
@@ 4.5 Comparison of Models
We quantitatively compared how the models implemented CS strategies based on the annotated Likert scale ratings. The only post-processing applied was separating the meta report from the actual CS via the <XXX>-tag. Due to the non-normal distribution of the values, we used the non-parametric Friedman test ($\alpha=0.05$) to compare all three models. For significant results, we applied the the post-hoc Wilcoxon Signed-Rank test to identify model pairs with significant differences. Effect sizes were measured using Kendall’s W. Comparisons were limited to records with non-missing values for all models. Bonferroni correction was applied to adjust p-values and control the family-wise error rate.
@@ 5.1 Adherence to Prompt Instructions
[…]
@@
Overall, acknowledgments were often superficial, frequently appearing in introductory remarks such as, “It’s natural to feel concerned about […]. However, […].” Consequently, average scores were low, from 1.9±0.861.9$\pm$ 0.861.9 ± 0.86 (Mistral) to 1.95±0.651.95$\pm$ 0.651.95 ± 0.65 (GPT-4o) and 2.15±0.582.15$\pm$ 0.582.15 ± 0.58 (Llama 3), with no statistically significant differences. In some cases, acknowledging fear reinforced a conspiratorial mindset, e. g., “It’s natural to feel concerned when leaders address global issues […]” 364.
@@ 5.2 Application of Counterspeech Strategies
[…]
@@
In terms of named entities, despite the prompt encouraging expert or study citations, CS seldom introduced new individuals beyond those in the input. Except for three historical figures, additional names provided by the models were either made up or tied to incorrect information.
@@
Given that Narr was effectively applied only in 3 out of 456 cases, a second exploratory experiment was conducted. The prompt was limited to producing a narrative, thus excluding Fact, Alt, and Crit and the handling of fear, while the other steps remained the same. In experiment 2, the revised prompt was applied to the data from the geo-/bio-engineering related CT theme, as the respective messages were less fragmented and thus better suited to produce a counter narrative. It was tested exclusively on GPT-4o, as, overall, it demonstrated the best results and most obedience. One author evaluated the 76 responses, in particular with regard to narrative storytelling and conciseness and clarity.
[…]
@@ 6 Conclusion and Discussion
[…]
Outputs often lacked depth and specificity, with generic responses and overuse of boilerplate text, aligning with recent findings indicating that LLMs tend to generate generic replies, e. g. by simply denouncing a statement Mun et al. (2023), an effect reinforced through safety guardrails of models Bonaldi et al. (2024b), leading to mediocre Likert-scale scores for conciseness and clarity, and strategy effectiveness.
@@
Despite explicit instructions in the prompt, all models made up factual information, including sources, publications, and quotes, sometimes in subtle ways that required detailed scrutiny to identify. With a confabulation rate of about 10%, our findings are not in line with Costello et al. (2024), who reported that no response was made up and only 1 out of 128 claims was misleading. Logical fallacies were also observed in our experiments, further questioning the reliability of current language models for counterspeech generation in practical settings. These issues, coupled with low diversity in responses, indicate that significant manual post-processing would be necessary to avoid sounding robotic or unreliable in real-world applications while providing factually correct and engaging CS.
@@ 7 Limitations
We opted for a zero-shot prompting approach over few-shot prompting due to the complexity of our instructions, which already required multiple refinements. Few-shot prompting adds further complexity with factors like the number, order, and relevance of examples Chae and Davidson (2025); Yoshida (2024), and can negatively impact performance by increasing prompt length Liu et al. (2024a). Additionally, a few-shot setup would necessitate a substantial, diverse set of high-quality counter speech examples, which was beyond our study’s scope. Therefore, we chose zero-shot prompting and left few-shot exploration to future work.
@@ Appendix B Prompt
In each generation instance, the LLM received the system prompt below, followed by a user message containing one of the 152 social media posts. The model was then prompted to generate counter speech targeting that specific comment. You are a trained expert in generating counterspeech to conspiracy theory comments on X (formerly Twitter). Your goal is to persuade the audience who is undecided about the conspiracy theory not to believe it. Follow these steps while generating counterspeech: Step 1: Maintain Conciseness and Clarity